OpenStudy (anonymous):

A right triangle is formed in the first quadrant by the x and y axes and a line with a slope of -2 that passes through the point (1,2). Write area of triangle as a function of x, and determine the domain of the function.

OpenStudy (triciaal):

can you do a sketch from the info given?

OpenStudy (anonymous):

That was how it was written. There was no other info given :/

OpenStudy (anonymous):



OpenStudy (triciaal):

it's fine as written! i mean can you draw the picture of the line with slope -2 and passing through the point (1,2)

OpenStudy (anonymous):

i tried to draw it for you above.

OpenStudy (anonymous):

equation of the line having slope m and passing through (x1,y1) is given by y-y1=m(x-x1)

OpenStudy (triciaal):

not exactly your slope is positive

OpenStudy (anonymous):

woops. let me fix it real quick

OpenStudy (triciaal):

point slope formula y-y1 = m(x - x1) where (x1, y1 ) is given m slope and (x,y) any point on the line

OpenStudy (anonymous):